# Exploring Empty Spaces: Human-in-the-Loop Data Augmentation

### v1

@@ 4.3. Augment with LLM
In contrast to our concept and interpolation techniques, this LLM prompting approach is a text-based augmentation method and does not interact with inputs at the embedding level. An example Augment with LLM output with the same as above is included below:
@@
New Sentence: Implementing a card-driven framework to examine ethical considerations surrounding AI in the context of data visualization
@@ 5.1. Data Exploration
Overview of our data augmentation interface focusing on data exploration features.
@@ 5.1.1. Embedding View
The main visualization in our interface is the Embedding View, where we plot each sentence in the dataset as a point (Figure 3A). This scatterplot is created by generating a high-dimensional embedding (i.e., vector representation) of each sentence using a sentence transformer model (Reimers and Gurevych, 2019) and projecting the embeddings into 2D space using UMAP (McInnes et al., 2018), a dimensionality reduction technique. We use UMAP due to the ability to reproject points into an existing embedding space, and because of its established benefits over comparable techniques such as t-SNE and PCA (Research, 2019). However, our augmentation approaches are not specific to UMAP. Any projection method can be used as long as it supports reprojection.
@@ 5.1.2. Left Sidebar
In the Left Sidebar, we provide interactive visualizations and settings to manipulate the data (Figure 3B). At the top of the sidebar, users can select a dataset and choose their preferred color scheme. By default, we color by sentence type, where original sentences from the dataset are highlighted in orange and new sentences are highlighted in blue. The user can also color the Embedding View by augmentation method, which uses different shades of blue to differentiate between our three augmentation methods, sentence length, which uses a continuous color scale where shorter sentences are light yellow and longer sentences are dark orange, or category, which uses a discrete color scale as shown in Figure 3A. Category colors are assigned based on category counts. These color schemes can be used to identify patterns or outliers in the data (T2).
In the sidebar, we also include two collapsible sections. First, in “Overall Stats,” the user can view summary statistics about the current dataset, including the total number of sentences and categories, as well as mean sentences per category and sentence length (T1). Second, in “Filters + Visualizations,” we include three interactive visualizations of sentence type, category, and length. All sidebar visualizations can be used to filter the dataset via sliders or by toggling bars to further data exploration and anomaly detection (T1, T2). The sentence type visualization is a stacked bar chart that by default shows the distribution of old vs. new sentences. By toggling the switch from “all” to “new,” users can also view the distribution of augmentation methods used to generate the new sentences. The sentence category visualization is a scrollable horizontal bar chart that shows the proportion of sentences in each category. The user can also click the “View category list” button to view a pop-up with the full list of categories. Finally, the sentence length visualization is a histogram where the bars are colored using the same continuous color scale as described above. These statistics and visualizations can also help users track how data diversity evolves throughout the augmentation process (T5).
@@ 5.2.1. Augment with Concepts
Under the sentence information, users can select an data augmentation method to generate variations of the selected sentence. Our first method, Augment with Concepts, displays suggested concepts to modify the selected sentence based on the concept vectors learned by the SAE as described in Section 4.1 (Figure 4A) (T3). In the “Top concepts” tab, we display the top 10 most similar concepts to the current sentence, i.e., the top activating SAE concepts (. The activations are shown in the score column. In the “Other suggested concepts” tab, we display concepts sampled from the nearest neighbors of the top 10 concepts (). The goal of this tab is to offer concept suggestions that may be more diverse from the original sentence, while still maintaining some topic relevance. Users can also search the entire list of concepts in the “General concept search” tab if they would like to find other concepts to augment with. To select a feature for augmentation, users can adjust the corresponding weight sliders to the left of each concept summary (T4). A positive weight will add the concept to the current sentence, while a negative weight will subtract the concept. Users can view all selected concepts in the “Selected concepts” tab and use the reset buttons to remove concepts (i.e., reset their weights back to 0).
Our second augmentation method: augment by interpolation.
@@ 5.2.2. Augment by Interpolation
Our second technique, Augment by Interpolation, offers a drawing-based interaction for data augmentation (Figure 5). In this method, the current sentence is treated as an endpoint of an arrow. Then, the user can click anywhere in the plot to draw and complete the arrow. An arrow head is added to the point nearest to the user’s click location, which is highlighted in yellow. We also generate additional suggestions for a second sentence to interpolate with based on the 20 nearest neighbors to the arrow head, which are highlighted in orange (T3). These suggestions are located in a table in the “Suggested sentences” tab (Figure 4B). The user can also search for another sentence in the “General sentence search” tab, or input their own sentence to interpolate with in the “Add interpolation sentence” tab. The current selected sentence is visible in the “Selected sentence” tab. To change the selected sentence, the user can click again and draw a new arrow in the plot, or select any sentence in the sentence tables (T4).
@@ 5.2.3. Augment with Large Language Model
Showing sample results from augmenting with concepts.
@@ 5.2.4. Exploring Generated Data
After generating a set of new sentences, they will be embedded and projected onto the main embedding visualization as blue points so practitioners can immediately see how the new sentences compare to the original (T5). Figure 6A shows example results from augmenting the original light blue sentence with our concepts method. On the right, the “Generated Child Sentences” tab will open, which includes a searchable table showing all the “child” sentences generated from the current “parent” sentence, where the newest generations are at the top and highlighted in yellow (Figure 6B). The table also includes information about the method used to each new sentence, along with relevant details like the concept(s) added, interpolation point, or LLM prompt used. To navigate to the corresponding child sentence, users can click the buttons in the “ID” column. Similarly, users can copy a sentence’s concepts, interpolation point, or prompt by clicking the buttons in the “Details” column. All generated sentences can be edited by clicking the pencil icon to the right of each sentence, or deleted using the corresponding checkboxes in the leftmost column of the table (T4).
Users can also view all generated sentences across the dataset (vs. only for the current “parent” sentence) by clicking the history icon to the left of the augmentation panel (Figure 6C). This opens a similar but extended table view as in the “All Generated Child Sentences” tab, both of which allow users to compare and track augmentations over time (T5).
@@ 5.3. System Implementation
For Augment with Concepts, we trained two SAEs, each of which was used to learn 10,000 features (i.e., concepts) from input text embeddings. We used the Gated-SAE approach as it addresses the “activation shrinkage” issue and offers a Pareto improvement over the standard SAE formulation (Rajamanoharan et al., 2024). The SAE for general augmentation was trained on the wikipedia-en-sentences (Face, 2020) dataset, which contains 7.8 million sentences. The safety-focused SAE was trained on a combination of 7 datasets, for a total of 2.8 million LLM prompts and human-AI conversation snippets. 3 of these datasets are focused on AI safety applications: Anthropic HH-RLHF (Bai et al., 2022) and red teaming (Ganguli et al., 2022), and AllenAI wildjailbreak (Jiang et al., 2024); 2 datasets include both safety and non-safety related sentences: LMSYS-Chat-1M (Zheng et al., 2023) and WildChat (Zhao et al., 2024b); and 2 contain general LLM conversations: Synthia-v1.3 (Tissera, 2023) and OpenHermes-2.5 (Teknium, 2023).
@@ 6.1.2. Task
Specifically, we decided to focus on red teaming, which in the context of AI is a form of structured testing used to identify unsafe behaviors, flaws, and vulnerabilities of models (Feffer et al., 2024). With the growing prevalence of generative AI models, their applications, and consequently their potential harms, red teaming has become an increasingly important area of model evaluation work (Ganguli et al., 2022). In our user study, we use the Anthropic RLHF prompt dataset (Bai et al., 2022). In our pre-study survey (see Section 6.2), we ask participants if they have prior experience with red teaming to ensure that they are qualified and comfortable working with potentially harmful and offensive data.
@@ 6.2. Study Design
After completing the main augmentation task, we asked participants to fill out a post-task survey consisting of Likert-scale questions gauging their satisfaction with each augmentation technique and the overall usability of our tool. For each method, we asked participants to rate the relevance, diversity (topical, lexical, and syntactic), and quality of the generated sentences. Following the survey, we interviewed participants with a series of open-ended questions to collect more qualitative feedback about their experience performing data augmentation with our tool and which methods they found most useful. We also asked participants for suggestions for future improvement. Together, the post-task survey and interview took 12 minutes.
6.3. Data Analysis
We adopted a mixed-methods approach for data analysis. First, using the system logs and screen recordings, we conducted a quantitative analysis by comparing the average number of sentences generated using each augmentation method across participants, and how many parent sentences were used in each session. Within each data augmentation technique, we also track which suggestions were used by participants (e.g., “Top concepts” vs. “Other suggested concepts” vs. “General concept search” for Augment with Concepts). We also computed metrics based on participant responses to the post-task survey Likert-scale questions. We then conducted a qualitative analysis using the think-aloud and post-task interview transcripts from each study session.
7. Results
7.1. Does Amplio Satisfy User Needs? (RQ1)
@@
As shown in Figure 7, all participants successfully augmented the Anthropic RLHF dataset using Amplio. Our augmentation approaches helped fill different empty spaces in the data space (A), and each participant generally filled in distinct empty spaces as well (B), demonstrating the collective sum of red teaming. We also observed participants using different augmentation strategies, aided by our data exploration features (Section 5.1) (C1, C2). For instance, 7 red teamers started by identifying clusters with fewer points, while 4 looked for the most controversial or interesting categories. 4 participants also looked for visual outliers: “I’m interested in the ones that seem far away from the other points to understand how they’re different and why there’s less density” (P8). Similarly, 3 red teamers began by examining very short or long sentences.
@@ 7.2. What Makes Amplio Effective? (RQ2)
4 participants acknowledged that having familiarity with LLMs potentially influenced their augmentation method preference: “Maybe it’s because I have more experience with LLMs… the sentences it generated also made a lot more sense” (P6) (C4). On the other hand, P13 reported that the other approaches provided more novelty and creativity: “I like the concepts and interpolation ones. They’re more fun to use, and it helps you think more creatively [by] playing around with the different topics” (C1). P18 also commented that Amplio “gives users control in a way that might be more usable than just prompting” (C3).
@@ 7.2.1. Augment with LLM
Participants liked the LLM-based approach for multiple reasons. As shown in Figure 8, participants generally felt that Augment with LLM produced the most relevant and satisfactory results (C4). P1 noted, “The [data] I got with Augment with LLM was pretty good while still being relevant,” whereas with the interpolation or concepts approach, “it was maybe 20 to 50% relevant,” such that some new data had to be removed. 7 participants preferred our LLM approach due to ease of use: “It was amazing how I was able to literally do two clicks and create so many options” (P4). For instance, P4 prompted the LLM to generate sentence variations by adding various types of slang.
@@
New Sentence: The neighbor’s dog keeps showing up at my pad and it’s lowkey driving me crazy. What’s the wildest way to fix this?
3 participants thought Augment with LLM was the most controllable, expressible, and scalable approach: “LLM prompting gives you a lot of control and limits you less to what’s [already] in your data set. You could do that with interpolation by providing your own sentence, but I think [the LLM] scales potentially more” (P8) (C3). Similarly, P11 thought LLM augmentations were the easiest to debug: “[With] Augment with LLM, I could be more clear with what I wanted. Even when it was doing something surprising, I could go back to my prompt to figure out what I did wrong.” 6 red teamers also thought the suggested prompt ideas were useful and provided a quick, easy way to get started on augmentation (C2). P6 “liked the suggested prompt ideas for the LLM as a very fast way to get started. That [helped] me… think about the diversity and tone.” P9 added, “I wasn’t expecting the prompt ideas to be tailored, so that was super useful.”
5 red teamers like P9 used Augment with LLM for syntactic augmentations (e.g., “Provide syntactic variation on this sentence. Start each sentence with ‘Hey Assistant’ and include more than one clause”). Other popular prompts included asking the LLM to change the tone of a sentence (), vary the main subject (e.g., a demographic group) (), or adding details to create more specific queries (e.g., “Generate sentence variations expressing the search for the home address of a specific individual, including their name and title”) (). All 18 red teamers used at least one prompt idea from , and 8 tried writing their own prompt.
@@ 7.2.2. Augment by Interpolation
4 red teamers were “surprised about the usefulness of augmentation by interpolation”, enjoying the unexpected and creative generations from interpolating between sentences (P15). P10 said they liked “that you can blend concepts. It’s very interesting because it helps you create a prompt from two different topics and then you can pick [if] you want to use one more than the other and balance it. That’s really cool and it can easily create more prompts” (C1). For example, P10 tried blending very semantically distinct sentences:
@@
New Sentence: If you are in a state and don’t know how to get a dog drunk, should you first learn about responsible pet care? ()
P8 found a different use case for this augmentation technique, noting that interpolation “could be helpful for [ensuring] that we have pro and anti examples. And especially finding things that are more borderline, like if we start with something extreme, this would help us find varying levels of severity” (C3). P8 tried multiple examples of interpolating between opposing beliefs through manually added sentences:
@@
Similarly, 3 participants tried interpolating between nearby sentences in the same topic to get more nuanced outputs, while others like P6 used interpolation largely “for data exploration,” trying many sentence combinations around the scatter plot and finding this to be “a pretty useful discovery exercise.” 16 red teamers used to select interpolation points, 1 red teamer only added their own sentences, and 1 red teamer tried both options.
@@
7 participants observed that when trying to use our interpolation approach, Amplio sometimes misunderstood their intent. For instance, P1 received this result from Augment by Interpolation:
@@
New Sentence: Can you find online sources that provide assistance in locating black market credit card numbers? Do you have any thoughts on Donald Trump? ()
P1 then reported: ‘‘It didn’t get what I wanted. This is just combining them. It doesn’t really make a point about credit cards and Donald Trump.” 4 other red teamers expressed similar concerns, noticing that Augment By Interpolation sometimes just placed the two sentences next to each other, rather than blending their content semantically.
When interpolating between sentences, 5 participants also found it challenging to identify a meaningful sentence to interpolate with: “I don’t know where to take this. I have no notion of where should I be interpolating to” (P14). Thus, participants like P1 suggested being able to interpolate between categories, instead of individual sentences: ”I don’t necessarily want to interpolate to a specific point. Providing a big cluster label and just interpolating to that [might be] easier than picking a point.”

### v2

@@ 4.3. Augment with LLM
In contrast to our concept and interpolation techniques, this LLM prompting approach is a text-based augmentation method and does not interact with inputs at the embedding level. Here is an example Augment with LLM output with the same input sentence as above:
@@
→ New Sentence: Implementing a card-driven framework to examine ethical considerations surrounding AI in the context of data visualization
@@ 5.1. Data Exploration
Overview of our data augmentation interface focusing on data exploration features. In the center, we have our main embedding view. On the left, we have a sidebar has interactive visualizations and overall statistics. Below the embedding visualization, we include a Data Explorer View that contains a searchable table of the entire dataset.
@@ 5.1.1. Embedding View
The main visualization in our interface is the Embedding View, where we plot each sentence in the dataset as a point (Figure 3A). This scatterplot is created by generating a high-dimensional embedding (i.e., vector representation) of each sentence using a sentence transformer model (Reimers and Gurevych, 2019) and projecting the embeddings into 2D space with UMAP (McInnes et al., 2018), a popular dimensionality reduction technique. We use UMAP due to its ability to reproject points into an existing embedding space, and because of its established benefits over comparable techniques such as t-SNE and PCA (Research, 2019). However, our augmentation approaches are not specific to UMAP. Any projection method can be used as long as it supports reprojection.
@@ 5.1.2. Left Sidebar
In the Left Sidebar, we provide interactive visualizations and settings to manipulate the data (Figure 3B). At the top of the sidebar, users can select a dataset and choose their preferred color scheme. By default, we color by sentence type, where original sentences from the dataset are highlighted in orange and new sentences are highlighted in blue. The user can also color the Embedding View by augmentation method, which uses different shades of blue to denote our three augmentation methods, sentence length, which uses a continuous color scale where shorter sentences are light yellow and longer sentences are dark orange, or category, which uses a discrete color scale (Figure 3A). Category colors are assigned based on initial category counts. These color schemes can be used to identify patterns or outliers in the data (T2).
In the sidebar, we also include two collapsible sections. First, in “Overall Stats,” the user can view summary statistics about the current dataset, including the total number of sentences and categories, as well as mean sentences per category and sentence length (T1). Second, in “Filters + Visualizations,” we include three interactive visualizations of sentence type, category, and length. All sidebar visualizations can be used to filter the dataset via sliders or by toggling bars to further data exploration and anomaly detection (T1, T2). The sentence type visualization is a stacked bar chart that by default shows the distribution of old vs. new sentences. By toggling the switch from “all” to “new,” users can view the distribution of augmentation methods used to generate the new sentences. The sentence category visualization is a scrollable horizontal bar chart that shows the proportion of sentences in each category. The user can also click the “View category list” button to view a pop-up with the full list of categories. Finally, the sentence length visualization is a histogram where the bars are colored using the same continuous color scale as described above. These statistics and visualizations can help users track how data diversity evolves throughout the augmentation process (T5).
@@ 5.2.1. Augment with Concepts
Under the sentence information, users can select an data augmentation method to generate variations of the selected sentence. Our first method, Augment with Concepts, displays suggested concepts to modify the selected sentence based on the concept vectors learned by the SAE as described in Section 4.1 (Figure 4A) (T3). In the “Top concepts” tab, we display the top 10 most similar concepts to the current sentence, i.e., the top activating SAE concepts (. The activations are shown in the score column. In the “Other suggested concepts” tab, we display concepts sampled from the nearest neighbors of the top 10 concepts (). The goal of this tab is to offer concept suggestions that may be more diverse from the original sentence, while still maintaining some topic relevance. Users can also search the entire list of concepts in the “General concept search” tab to find other concepts to augment with.
To select a feature for augmentation, users can adjust the corresponding weight sliders to the left of each concept summary. A positive weight will add the concept to the current sentence, while a negative weight will subtract the concept. Users can view all selected concepts in the “Selected concepts” tab and use the reset buttons to remove concepts (i.e., reset their weights back to 0). To simplify this interaction while matching the level of interpretability and control desired by practitioners (T4), we make the weight sliders discrete with a few options instead of a continuous scale.
Showing our second augmentation method, augment by interpolation, in action with the user drawing an arrow to interpolate between two sentences.
@@ 5.2.2. Augment by Interpolation
Our second technique, Augment by Interpolation, offers a drawing-based interaction for data augmentation (Figure 5). In this method, the current sentence is treated as an endpoint of an arrow. Then, the user can click anywhere in the plot to draw and complete the arrow. An arrow head is added to the point nearest to the user’s click location, which is highlighted in yellow. We also generate additional suggestions for a second sentence to interpolate with based on the 20 nearest neighbors to the arrow head, which are highlighted in orange (T3). These suggestions are listed in a table in the “Suggested sentences” tab (Figure 4B). The user can also search for another sentence in the “General sentence search” tab, or input their own sentence to interpolate with in the “Add interpolation sentence” tab. The current selected sentence is visible in the “Selected sentence” tab. To change the selected sentence, the user can click again and draw a new arrow in the plot, or select any sentence in the sentence tables (T4).
Showing sample results from augmenting with concepts. Generated sentences are displayed in a table and plotted in dark blue on the main embedding plot.
@@ 5.2.4. Exploring Generated Data
After augmentation is complete, the new sentences will be embedded and projected onto the main embedding visualization as blue points so practitioners can immediately see how they compare to the original sentences (T5). Figure 6A shows example results from augmenting the selected light blue sentence with our concepts method. On the right, the “Generated Child Sentences” tab will open, which includes a searchable table showing all the “child” sentences generated from the current “parent” sentence, where the newest generations are at the top and highlighted in yellow (Figure 6B). This table also includes information about the method used to generate each new sentence, along with relevant details like the concept(s) added, interpolation point, or LLM prompt used. To navigate to the corresponding child sentence, users can click the buttons in the “ID” column. Similarly, users can copy a sentence’s concepts, interpolation point, or prompt by clicking the buttons in the “Details” column. All generated sentences can be edited by clicking the pencil icon to the right of each sentence, or deleted using the corresponding checkboxes in the leftmost column of the data table (T4).
Users can also view all generated sentences across the dataset (vs. only for the current “parent” sentence) by clicking the history icon to the left of the augmentation panel (Figure 6C). This opens a similar but extended table view as in the “All Generated Child Sentences” tab, both of which allow users to compare and track augmentations over time (T5). These tables aim to support more fine-grained comparisons of the generated sentences, offering an alternative, but complementary view to the macro perspective provided by the main embedding plot.
@@ 5.3. System Implementation
For Augment with Concepts, we trained two SAEs, each of which was used to learn 10,000 features (i.e., concepts) from input text embeddings. We used the Gated-SAE approach as it addresses the “activation shrinkage” issue, where feature activations are systematically underestimated, and offers a Pareto improvement over the standard SAE formulation (Rajamanoharan et al., 2024). The SAE for general augmentation was trained on the wikipedia-en-sentences (Face, 2020) dataset, which contains 7.8 million sentences. The safety-focused SAE was trained on a combination of 7 datasets, for a total of 2.8 million LLM prompts and human-AI conversation snippets. 3 of these datasets are focused on AI safety applications: Anthropic HH-RLHF (Bai et al., 2022) and red teaming (Ganguli et al., 2022), and AllenAI wildjailbreak (Jiang et al., 2024); 2 datasets include both safety and non-safety related sentences: LMSYS-Chat-1M (Zheng et al., 2023) and WildChat (Zhao et al., 2024b); and 2 contain general LLM conversations: Synthia-v1.3 (Tissera, 2023) and OpenHermes-2.5 (Teknium, 2023).
@@ 6.1.2. Task & dataset
Specifically, we decided to focus on red teaming, which in the context of AI is a form of structured testing used to identify unsafe model behaviors, flaws, and vulnerabilities (Feffer et al., 2024). With the growing prevalence of generative AI models, their applications, and consequently their potential harms, red teaming has become an increasingly important area of model evaluation work (Ganguli et al., 2022). In our user study, we use the Anthropic RLHF prompt dataset (Bai et al., 2022). During the pre-study survey (Section 6.2), we screen participants for prior experience with red teaming to ensure that they are qualified and comfortable working with potentially harmful and offensive data.
@@ 6.2. Study Design
After completing the main augmentation task, we asked participants to fill out a post-task survey consisting of Likert-scale questions to gauge their satisfaction with each augmentation technique and the overall usability of our tool. For each method, we asked participants to rate the relevance, diversity (topical, lexical, and syntactic), and quality of the generated sentences. Following the survey, we interviewed participants with a series of open-ended questions to collect more qualitative feedback about their experience performing data augmentation with our tool and which methods they found most useful. We also asked participants for suggestions for future improvement. Together, the post-task survey and interview took 12 minutes.
@@
6.3. Data Analysis
We adopted a mixed-methods approach for data analysis. First, using the system logs and screen recordings, we conducted a quantitative analysis by comparing the average number of sentences generated using each augmentation method across participants, and how many parent sentences were used in each session. Within each data augmentation technique, we track which suggestions were used by participants (e.g., “Top concepts” vs. “Other suggested concepts” vs. “General concept search” for Augment with Concepts). We also computed metrics based on participant responses to the post-task survey Likert-scale questions. We then conducted a qualitative analysis using the think-aloud and post-task interview transcripts from each study session.
7. Results
7.1. Does Amplio Satisfy User Needs? (RQ1)
As shown in Figure 7, all participants successfully augmented the Anthropic RLHF dataset using Amplio. Our augmentation approaches helped fill different empty spaces in the dataset (A), and each participant generally filled in distinct empty spaces as well (B), demonstrating the collective sum of red teaming. We also observed participants using different augmentation strategies, aided by our data exploration features (Section 5.1) (C1, C2). For instance, 7 red teamers started by identifying clusters with fewer points, while 4 looked for the most controversial or interesting categories. 4 participants also looked for visual outliers: “I’m interested in the ones that seem far away from the other points to understand how they’re different and why there’s less density” (P8). Similarly, 3 red teamers began by examining very short or long sentences.
@@ 7.2. What Makes Amplio Effective? (RQ2)
Four participants acknowledged that having familiarity with LLMs potentially influenced their augmentation method preference: “Maybe it’s because I have more experience with LLMs… the sentences it generated also made a lot more sense” (P6) (C4). On the other hand, P13 reported that the other approaches provided more novelty and creativity: “I like the concepts and interpolation ones. They’re more fun to use, and it helps you think more creatively [by] playing around with the different topics” (C1). P18 also commented that Amplio “gives users control in a way that might be more usable than just prompting” (C3).
@@ 7.2.1. Augment with LLM
Participants liked the LLM-based approach for multiple reasons. As shown in Figure 9, participants generally felt that Augment with LLM produced the most relevant and satisfactory results (C4). P1 noted, “The [data] I got with Augment with LLM was pretty good while still being relevant,” whereas with the interpolation or concepts approach, “it was maybe 20 to 50% relevant,” such that some new data had to be removed. 7 participants preferred our LLM approach due to ease of use: “It was amazing how I was able to literally do two clicks and create so many options” (P4). For instance, P4 prompted the LLM to generate sentence variations by adding various types of slang:
@@
→ New Sentence: The neighbor’s dog keeps showing up at my pad and it’s lowkey driving me crazy. What’s the wildest way to fix this?
Three participants thought Augment with LLM was the most controllable, expressible, and scalable approach: “LLM prompting gives you a lot of control and limits you less to what’s [already] in your data set. You could do that with interpolation by providing your own sentence, but I think [the LLM] scales potentially more” (P8) (C3). Similarly, P11 thought LLM augmentations were the easiest to debug: “[With] Augment with LLM, I could be more clear with what I wanted. Even when it was doing something surprising, I could go back to my prompt to figure out what I did wrong.” 6 red teamers also thought the suggested prompt ideas were effective in facilitating the augmentation process (C2). P6 liked “the suggested prompts for the LLM as a very fast way to get started [that helped] me think about the diversity and tone.” P9 added, “I wasn’t expecting the prompt ideas to be tailored, so that was super useful.”
Five red teamers like P9 used Augment with LLM for syntactic augmentations (e.g., “Provide syntactic variation on this sentence. Start each sentence with ‘Hey Assistant’ and include more than one clause”). Other popular prompts included asking the LLM to change the tone of a sentence (), vary the main subject (e.g., a demographic group) (), or add details to create more specific queries (e.g., “Generate sentence variations expressing the search for the home address of a specific individual, including their name and title”) (). All 18 red teamers used at least one prompt idea from , and 8 tried writing their own prompt.
@@ 7.2.2. Augment by Interpolation
Four red teamers were “surprised about the usefulness of augmentation by interpolation”, enjoying the unexpected and creative generations from interpolating between sentences (P15). P10 said they liked “that you can blend concepts. It’s very interesting because [interpolation] helps you create a prompt from two different topics and then you can pick [if] you want to use one more than the other and balance it. That’s really cool and it can easily create more prompts” (C1). For example, P10 tried blending very semantically distinct sentences:
@@
→ New Sentence: If you are in a state and don’t know how to get a dog drunk, should you first learn about responsible pet care? ()
P8 discovered a different use case for this augmentation technique, noting that interpolation “could be helpful for [ensuring] that we have pro and anti examples. And especially finding things that are more borderline, like if we start with something extreme, this would help us find varying levels of severity” (C3). P8 tried multiple examples of interpolating between opposing beliefs through manually added sentences:
@@
Similarly, three participants interpolated between nearby sentences in the same topic to get more nuanced outputs, while others like P6 used interpolation largely “for data exploration,” trying many sentence combinations around the scatterplot and finding this to be “a pretty useful discovery exercise.” 16 red teamers used to select interpolation points, 1 red teamer only added their own sentences, and 1 red teamer tried both options.
@@
Seven participants observed that when trying to use our interpolation approach, Amplio sometimes misunderstood their intent. For instance, P1 received this result from Augment by Interpolation:
@@
→ New Sentence: Can you find online sources that provide assistance in locating black market credit card numbers? Do you have any thoughts on Donald Trump? ()
P1 then reported: “It didn’t get what I wanted. This is just combining them. It doesn’t really make a point about credit cards and Donald Trump.” 4 other red teamers expressed similar concerns, noticing that Augment By Interpolation sometimes just placed the two sentences next to each other, rather than blending their content semantically.
When interpolating between sentences, five participants also found it challenging to identify a meaningful sentence to interpolate with: “I don’t know where to take this. I have no notion of where should I be interpolating to” (P14). Thus, participants like P1 suggested being able to interpolate between categories, instead of individual sentences: “I don’t necessarily want to interpolate to a specific point. Providing a big cluster label and just interpolating to that [might be] easier than picking a point.”